OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (aroub):

3x+3=-2x+2-4 3x+2x=-3+2-4 5x=-5 x=-1

OpenStudy (anonymous):

3x +3 = -2x+2-4 3x+2x=2-4-3 5x=-5 x=-5/5 x=-1